A colourless compound weighing 0.84 on heating gives 0.22 g of CO2(carbon dioxide) and 0.09 g of H2O(water) and leaves a white residue.What is the molecular mass of the compound? Pls give the steps
1
Expert's answer
2013-12-02T07:13:32-0500
Calculate the amount of the released CO2 and H2O:

n(CO2) = m/M = 0.22/44 = 0.005 mol;
n(H2O) = m/M = 0.09/18 = 0.005 mol;

The amount of substance of carbon dioxide is equal to the water amount, therefore we can assume that the amount of substance of initial compound is same. No we can calculate the molar mass:

M = m/n = 0.84/0.005 = 168 g/mol.